OpenStudy (anonymous):

imagine you are teaching a fellow student how to solve: 3 − 5(x + 2) = 3(x + 1) + 6 In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (gorv):

multiply -5 and 3 inside the bracket and open it 3-5*x-5*2=3x+3+6 3-5x-10=3x+9 -5x-7=3x+9 -5x-3x=9+7 -8x=16 x=-2

OpenStudy (gorv):

@Sosa_ got it or not???